# Barron Space Representations for Elliptic PDEs with Homogeneous Boundary Conditions

### v1

@@ 3.2.2 Analysis of the Barron Norm
By a similar argument of absolute and uniform convergence as in the proof of (3.25), the following computation is justified:
$A_{ij}\partial_{ij}u$ $=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\quad\quad\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{m=1}^{d}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)\pm\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$=\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}S_{v_{n}}(x),$
where each $v_{n}=\left(\pm k_{1}\pm k_{1}^{\prime},\ldots,\pm k_{d}\pm k_{d}^{\prime}\right)$, with appropriately chosen signs such that $v_{n}\in\mathbb{N}^{d}$.
Combining the reordering invariance of the Barron norm (Remark 2.9), a similar summation–differentiation interchange argument in the proof of Proposition 3.11 (4), and Proposition 3.10, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ij}\partial_{ij}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$ $=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}(I-\Delta)^{-1}_{D}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{n=1}^{2^{d}}\pm\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\frac{1}{1+\pi^{2}\|v_{n}\|^{2}}S_{v_{n}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
@@
    Suppose the condition on $c(x)$ in Assumption 2.11 holds, and let $u\in\mathcal{B}_{s}^{2}(\Omega)$. Then we have
    $\|(I-\Delta)^{-1}_{D}(cu)\|_{{\mathcal{B}}_{s}^{2}(\Omega)}\leq\ell_{c}\|u\|_{{\mathcal{B}}_{s}^{2}(\Omega)}.$
@@
    Suppose the condition on $c(x)$ in Assumption 2.12 holds, and let $u\in{\mathcal{B}}_{c}^{2}(\Omega)$. Then we have
    $\|(I-\Delta)^{-1}_{N}(cu)\|_{{\mathcal{B}}_{c}^{2}(\Omega)}\leq\ell_{c}\|u\|_{{\mathcal{B}}_{c}^{2}(\Omega)}.$
@@
The result follows from the linearity of $(I-\Delta)^{-1}_{D}$ and $(I-\Delta)^{-1}_{N}$, together with (3.22), (3.23), Proposition 3.11, and Lemmas 3.12 and 3.13. ∎
@@ 3.3 Step 3: Neural Network Approximation
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
If we regard $\Theta$ as a random variable on the probability space $\left((\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k},{\mu(a,w,b)}^{\otimes k}\right)$, then to prove the theorem, it suffices to establish that
@@
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$ $=\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\int_{\Omega}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}{\mu(a,w,b)}^{\otimes k}$
@@
$\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\langle w_{i},e_{j}\rangle\sigma^{\prime}(w_{i}^{\top}x+b_{i})-\partial_{j}g(x)\right)\left(a_{\ell}\langle w_{\ell},e_{j}\rangle\sigma^{\prime}(w_{\ell}^{\top}x+b_{\ell})-\partial_{j}g(x)\right){\,\mathrm{d}}\mu^{\otimes k}=0,$
@@
$\mathbb{E}_{\mu^{\otimes k}}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$ $=\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\int_{\Omega}\left(\partial_{j}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}\mu^{\otimes k}$
@@
$\mathbb{E}_{\mu^{\otimes k}}\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2}$ $=\mathbb{E}_{\mu^{\otimes k}}\sum_{j=1}^{d}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$
$\leq\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right)^{2}\right]{\,\mathrm{d}}x$
$\leq\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\right)^{2}\right]{\,\mathrm{d}}x=\frac{1}{k}\int_{\Omega}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]{\,\mathrm{d}}x$
$\leq\frac{m(\Omega)}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]=\frac{1}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right].$ (3.37)
@@ Appendix A Omitted Proofs in Sections 2 and 3
[…]
@@
By construction, $\widetilde{g}$ is odd in each coordinate, and hence its expansion in the basis (A.1) contains only sine terms:
@@
[…]
@@
By construction, $\widetilde{g}$ is even in each coordinate, and hence its expansion in the basis (A.1) contains only cosine terms:
[…]
The remaining part of the proof follows the same argument as in (2).
[…]
@@
By construction, $\widetilde{g}$ is odd in the $i$-th and $j$-th coordinates and even in the remaining ones. As a result, its expansion in the basis (A.1) consists of sine terms in the $i$-th and $j$-th positions and cosine terms in the remaining positions:
[…]
The remaining part of the proof follows the same argument as in (2).
Example A.1.
We present an example demonstrating that a function in $L^{2}(\Omega)$ may not have a unique expansion in terms of the exponential basis (1).
Consider the constant function $g(x)=1$ on $\Omega$. Then $g$ admits at least two distinct expansions with respect to the exponential basis:
$g(x)=e^{i\pi 0^{\top}x}=\sum_{\omega\in\mathbb{Z}^{d}}a_{\omega}e^{i\pi\omega^{\top}x}\quad\text{in }L^{2}(\Omega),$
where one possible choice of the coefficients $\{a_{\omega}\}$ is given by
$a_{\omega}=\begin{cases}\displaystyle\frac{1}{2^{d-1}}-1,&\omega=0,\\ \displaystyle\frac{1}{2^{d-1}}\prod_{1\leq j\leq d,\,\omega_{j}\neq 0}\frac{1-e^{-i\pi\omega_{j}}}{i\pi\omega_{j}},&\text{otherwise}.\end{cases}$
[…]
A.2 Proof of Lemma 3.6
Let $A(x)=O(x)\Lambda(x)O(x)^{\top}$ be the orthogonal decomposition of $A(x)$, where $O(x)$ is an orthogonal matrix and $\Lambda(x)=\mathrm{diag}\{\lambda_{1}(x),\dots,\lambda_{d}(x)\}$ is a diagonal matrix whose diagonal entries are the eigenvalues of $A(x)$. We claim that
$a_{\max}=\sup_{x\in\Omega}\sup_{1\leq i\leq d}|\lambda_{i}(x)|.$
Indeed, since orthogonal transformations preserve the Euclidean norm, we have
$a_{\max}$ $=\|A(x)\|_{\mathrm{op}}=\sup_{x\in\Omega}\sup_{v\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|A(x)v\|}{\|v\|}=\sup_{x\in\Omega}\sup_{v\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|O(x)\Lambda(x)O(x)^{\top}v\|}{\|O(x)^{\top}v\|}$
$=\sup_{x\in\Omega}\sup_{\xi\in\mathbb{R}^{d}\setminus\{0\}}\frac{\|\Lambda(x)\xi\|}{\|\xi\|}=\sup_{x\in\Omega}\sup_{\xi\in\mathbb{R}^{d}\setminus\{0\}}\left(\frac{\lambda_{1}(x)^{2}\xi_{1}^{2}+\cdots+\lambda_{d}(x)^{2}\xi_{d}^{2}}{\xi_{1}^{2}+\cdots+\xi_{d}^{2}}\right)^{1/2}$
$=\sup_{x\in\Omega}\sup_{1\leq i\leq d}|\lambda_{i}(x)|.$
Therefore, by the Cauchy–Schwarz inequality, we obtain
$\int_{\Omega}\nabla u\cdot A\nabla v{\,\mathrm{d}}x$ $=\int_{\Omega}(\nabla uO(x))\cdot\Lambda(x)\cdot(\nabla vO(x))^{\top}{\,\mathrm{d}}x$
$\leq a_{\max}\|\nabla uO(x)\|_{L^{2}(\Omega)}\|\nabla vO(x)\|_{L^{2}(\Omega)}$
$=a_{\max}\|\nabla u\|_{L^{2}(\Omega)}\|\nabla v\|_{L^{2}(\Omega)}.$
A.3 The Poincaré Constant for the Unit Hypercube
@@
$\left\{S_{k}(x)=\prod_{i=1}^{d}\sin(\pi k_{i}x_{i}):k=(k_{1},\ldots,k_{d})\in\mathbb{N}^{d},\ k_{i}\neq 0\text{ for all }i,\ x=(x_{1},\ldots,x_{d})\in\Omega\right\},$
@@
$|g^{\prime}(z)-g_{m}^{\prime}(z)|=\left|g^{\prime}(z)-\frac{g(z_{j+1})-g(z_{j})}{h_{1}}\right|=|g^{\prime}(z)-g^{\prime}(\xi)|=|g^{\prime\prime}(\eta)||z-\xi|\leq|g^{\prime\prime}(\eta)|h_{1}\leq\frac{2\sqrt{d}B}{m}.$
@@
$\|g-g_{m}\|_{H^{1}(I_{d})}^{2}=\|g-g_{m}\|_{L^{2}(I_{d})}^{2}+\|g^{\prime}-g_{m}^{\prime}\|_{L^{2}(I_{d})}^{2}\leq 2\sqrt{d}\|g-g_{m}\|_{L^{\infty}(I_{d})}^{2}+2\sqrt{d}\|g^{\prime}-g_{m}^{\prime}\|_{L^{\infty}(I_{d})}^{2}.$
@@
$\mathscr{G}\coloneqq\left\{c+A\operatorname{ReLU}(w^{\top}x+b):|c|\leq 2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ |A|\leq 8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ \|w\|=1,\ |b|\leq\sqrt{d}\right\}.$
@@
Acknowledgements
The work of ZC is supported in part by the National Science Foundation via grant DMS-2509011. ZC thanks Jianfeng Lu and Yulong Lu for helpful discussions. LH gratefully acknowledges Padmavathi Srinivasan for her continuous encouragement and patient guidance during his first two years of graduate study. He also thanks Wenkui Du and Shengxuan Zhou for many valuable discussions.

### v2

@@ 3.2.2 Analysis of the Barron Norm
We compute
$A_{ij}\partial_{ij}u$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\left(\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\cos\left(\pi k_{m}x_{m}\right)\sin\left(\pi k_{m}^{\prime}x_{m}\right)\right)$
$\qquad\cdot\sin\left(\pi k_{i}x_{i}\right)\cos\left(\pi k_{i}^{\prime}x_{i}\right)\sin\left(\pi k_{j}x_{j}\right)\cos\left(\pi k_{j}^{\prime}x_{j}\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\pi^{2}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\cdot\frac{1}{2^{d}}\prod_{\begin{subarray}{c}1\leq m\leq d\\ m\neq i,j\end{subarray}}\left(\sin\left(\pi(k_{m}+k_{m}^{\prime})x_{m}\right)-\sin\left(\pi(k_{m}-k_{m}^{\prime})x_{m}\right)\right)$
$\quad\cdot\left(\sin(\pi(k_{i}+k_{i}^{\prime})x_{i})+\sin(\pi(k_{i}-k_{i}^{\prime})x_{i})\right)\left(\sin(\pi(k_{j}+k_{j}^{\prime})x_{j})+\sin(\pi(k_{j}-k_{j}^{\prime})x_{j})\right)$
$=$ $\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\epsilon_{i}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot S_{k+\epsilon\circ k^{\prime}}(x),$
Thus, we obtain
$\|(I-\Delta)^{-1}_{D}(A_{ij}\partial_{ij}u)\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\epsilon_{i}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot(I-\Delta)^{-1}_{D}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
$=$ $\left\|\sum_{k\in\mathbb{N}^{d}}\sum_{k^{\prime}\in\mathbb{N}^{d}}\sum_{\epsilon\in\{\pm 1\}^{d}}\frac{\pi^{2}}{2^{d}}a_{k}b_{k^{\prime}}k_{i}^{\prime}k_{j}^{\prime}\epsilon_{i}\epsilon_{j}\prod_{m=1}^{d}\epsilon_{m}\cdot\frac{1}{1+\pi^{2}\|k+\epsilon\circ k^{\prime}\|^{2}}S_{k+\epsilon\circ k^{\prime}}\right\|_{\mathcal{B}_{s}^{2}(\Omega)}$
@@
    Suppose the condition on $c(x)$ in Assumption 2.5 holds, and let $u\in\mathcal{B}_{s}^{2}(\Omega)$. Then we have (I-Δ)^-1_D(cu)_B_s^2(Ω)≤ℓ_cu_B_s^2(Ω).
@@
    Suppose the condition on $c(x)$ in Assumption 2.6 holds, and let $u\in{\mathcal{B}}_{c}^{2}(\Omega)$. Then we have (I-Δ)^-1_N(cu)_B_c^2(Ω)≤ℓ_cu_B_c^2(Ω).
@@
The result follows from the linearity of $(I-\Delta)^{-1}_{D}$ and $(I-\Delta)^{-1}_{N}$, together with (3.20), (3.21), Proposition 3.7, and Lemmas 3.8 and 3.9. ∎
@@ 3.3 Step 3: Neural Network Approximation
$\theta_{\omega}\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\begin{cases}\arctan\left(c_{\omega,I}/c_{\omega,R}\right),&\text{if }c_{\omega,R}\neq 0,\\ \operatorname{sgn}(c_{\omega,I}){\pi}/2,&\text{if }c_{\omega,R}=0.\end{cases}$
@@
$Z\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\sum_{\omega\in\mathbb{Z}^{d}}|c_{\omega}|\leq\|g\|_{\mathcal{B}_{e}^{2}(\Omega)}+\varepsilon<\infty.$
@@
$\mu(a,w,b)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\sum_{\omega\in\mathbb{Z}^{d}}\frac{|c_{\omega}|}{Z}\delta_{(Z,\pi\omega,\theta_{\omega})}$
@@
$\Theta\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\left\{(a_{i},w_{i},b_{i})\right\}_{1\leq i\leq k}\in(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k},$
@@
$R(\Theta;x)\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)-g(x),$
which is a random variable on the probability space $\left((\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k},{\mu(a,w,b)}^{\otimes k}\right)$. Then to prove the theorem, it suffices to establish that
@@
$\mathbb{E}_{\mu^{\otimes k}}\left\|R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$
$=$ $\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\int_{\Omega}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma(w_{i}^{\top}x+b_{i})-g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}{\mu(a,w,b)}^{\otimes k}$
@@
$\int_{(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R})^{k}}\left(a_{i}\langle w_{i},e_{j}\rangle\sigma^{\prime}(w_{i}^{\top}x+b_{i})-\partial_{j}g(x)\right)$
$\qquad\qquad\qquad\qquad\cdot\left(a_{\ell}\langle w_{\ell},e_{j}\rangle\sigma^{\prime}(w_{\ell}^{\top}x+b_{\ell})-\partial_{j}g(x)\right){\,\mathrm{d}}\mu^{\otimes k}=0,$
and
@@
$\mathbb{E}_{\mu^{\otimes k}}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}$
$=$ $\int_{\left(\mathbb{R}\times\mathbb{R}^{d}\times\mathbb{R}\right)^{k}}\int_{\Omega}\left(\partial_{j}\left(\frac{1}{k}\sum_{i=1}^{k}a_{i}\sigma\left(w_{i}^{\top}x+b_{i}\right)\right)-\partial_{j}g(x)\right)^{2}{\,\mathrm{d}}x\mathrm{d}\mu^{\otimes k}$
@@
$\begin{split}&\mathbb{E}_{\mu^{\otimes k}}\|\nabla R(\Theta;x)\|_{L^{2}(\Omega)}^{2}=\mathbb{E}_{\mu^{\otimes k}}\sum_{j=1}^{d}\left\|\partial_{j}R(\Theta;x)\right\|^{2}_{L^{2}(\Omega)}\\ \leq&\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\sigma^{\prime}\left(w^{\top}x+b\right)\right)^{2}\right]{\,\mathrm{d}}x\\ \leq&\frac{1}{k}\int_{\Omega}\sum_{j=1}^{d}\mathbb{E}_{\mu(a,w,b)}\left[\left(a\left\langle w,e_{j}\right\rangle\right)^{2}\right]{\,\mathrm{d}}x=\frac{1}{k}\int_{\Omega}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]{\,\mathrm{d}}x\\ \leq&\frac{m(\Omega)}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right]=\frac{1}{k}\mathbb{E}_{\mu(a,w,b)}\left[a^{2}\left\|w\right\|^{2}\right].\end{split}$
@@ Appendix A Omitted Proofs in Sections 2 and 3
Given a function $g\in L^{2}(\Omega)$, we define its coordinatewise odd extension by
@@
Since $\widetilde{g}$ is odd in each coordinate, its expansion in (A.1) contains only sine terms:
@@
(3). The proof follows the same argument as in (2) by considering the coordinatewise even extension of $g\in L^{2}(\Omega)$, say
@@
(4). The proof follows the same argument as in (2) by considering an extension of $g\in L^{2}(\Omega)$ that is odd in the $i$-th and $j$-th coordinates and even in the remaining ones, say
@@
A.2 The Poincaré Constant for the Unit Hypercube
@@
$\left\{S_{k}(x)=\prod_{i=1}^{d}\sin(\pi k_{i}x_{i}):k=(k_{1},\ldots,k_{d})\in\mathbb{N}_{+}^{d}\right\},$
@@
$|g^{\prime}(z)-g_{m}^{\prime}(z)|$ $=\left|g^{\prime}(z)-\frac{g(z_{j+1})-g(z_{j})}{h_{1}}\right|=|g^{\prime}(z)-g^{\prime}(\xi)|$
$=|g^{\prime\prime}(\eta)||z-\xi|\leq|g^{\prime\prime}(\eta)|h_{1}\leq\frac{2\sqrt{d}B}{m}.$
@@
$\|g-g_{m}\|_{H^{1}(I_{d})}^{2}$ $=\|g-g_{m}\|_{L^{2}(I_{d})}^{2}+\|g^{\prime}-g_{m}^{\prime}\|_{L^{2}(I_{d})}^{2}$
$\leq 2\sqrt{d}\|g-g_{m}\|_{L^{\infty}(I_{d})}^{2}+2\sqrt{d}\|g^{\prime}-g_{m}^{\prime}\|_{L^{\infty}(I_{d})}^{2}.$
@@
$\mathscr{G}\coloneqq\Big\{c+A\operatorname{ReLU}(w^{\top}x+b):$
$\qquad\qquad\qquad\qquad|c|\leq 2\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ |A|\leq 8\sqrt{d}\|g\|_{{\mathcal{B}}_{s}^{2}(\Omega)},\ \|w\|=1,\ |b|\leq\sqrt{d}\Big\}.$
@@
Acknowledgments
Z. Chen thanks Jianfeng Lu and Yulong Lu for helpful discussions. L. Huang gratefully acknowledges Padmavathi Srinivasan for her continuous encouragement and patient guidance during his first two years of graduate study. He also thanks Wenkui Du, Lev Fedorov, Mengxuan Yang, and Shengxuan Zhou for many valuable discussions. We used ChatGPT for minor language editing; all technical content was written and verified by the authors.